3) An airplane has 2 engines. P(malfunction) = .0004 for a single engine on a single flight and these malfunctions are independent. A flight will crash if both engines malfunction on that flight. The plane makes 3 roundtrips per day [which is 6 flights]. Find:
[hint: numbers like 4.95E-5 should be written as .0000495]

a) P(first engine use does not malfunction) =

b) P(first flight has no engine malfunctions) =

c) P(no engine malfunctions in first week) =

d) P(at least one engine malfunction in first year) =
1
Expert's answer
2020-12-14T13:09:06-0500

a) P(first engine use does not malfunction) = 1 - P(malfunction) = 1-0.0004 = 0.9996

b) P(first flight has no engine malfunctions) = (1 - P(malfunction))2 = 0.99962 = 0.9992

c) P(no engine malfunctions in first week) = (P(first flight has no engine malfunctions))6*7 = 0.999242 = 0.966945

d) P(at least one engine malfunction in first year) = 1-(P(first flight has no engine malfunctions))6*365 = 1-0.99922190 = 0.826695